OpenStudy (anonymous):

What is the slope of the line which passes through (–6, 0) and (4, 0)? What is the slope of the line which passes through (–6, 0) and (4, 0)? @Mathematics

OpenStudy (diyadiya):

slope=y2-y1/x2-x1 =0/10=0

OpenStudy (anonymous):

its a horizontal line

OpenStudy (anonymous):

slope = y^2(y2) - y^1(y1) / x2 - x1 = 0/10=0 :)

OpenStudy (anonymous):

thanks